OpenStudy (anonymous):

Lauren is pulling on a trunk with a 15-N force. Wayne pulls at right angles to Lauren. With how much force must Wayne pull to make the resultant force on the trunk 22 N? Round answer to one decimal place.

OpenStudy (anonymous):

options are: . 26.6 N B. 21.4 N C. 18.7 N D. 16.1 N E. 17.4 N

OpenStudy (mathmate):

forces are at 90 degrees, so resultant can be calculated by Pythagoras theorem. L=15 W=Wayne's pulling force Resultant = sqrt(L^2+W^2)=22 Solve to get W=sqrt(22^2-15^2)

OpenStudy (anonymous):

so its 16.1?

OpenStudy (mathmate):

Correct!

OpenStudy (anonymous):

Thank you!

OpenStudy (mathmate):

You're welcome!